OpenStudy (anonymous):

Which is a solution of the equation? 2x + y = 5 5 (2, –1) (1, 2) (3, –1) What is the slope of the line that passes through the points (3, –1) and (–2, –5)? A. 5/4 B. - 5/4 C. 4/5 D. -4/5

OpenStudy (anonymous):

plug in the values

OpenStudy (anonymous):

for example 2(3)-1=???

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

which one did you get?

OpenStudy (anonymous):

so is the first one (3, -1)

OpenStudy (anonymous):

2(3)-1= 5 correct

OpenStudy (haseeb96):

m= y2-y1/x2-x1 m is the slope use this formula and get the slope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the second one b

OpenStudy (haseeb96):

(3,-1) is the correct answer and D is the correct answer for slope

OpenStudy (haseeb96):

sorry C is the correct answer for slope

OpenStudy (haseeb96):

m=-5+1/-2-3 =-4/-5 =4/5

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can y'all help me with more

OpenStudy (anonymous):

With pleasure my listener.

OpenStudy (anonymous):

What is the slope of the line passing through points A and B? A. -2 B. -1/2 C.1/2 D.2

OpenStudy (anonymous):

First, we must write the equation of a line given the slope and the coordinates of one point on it. For example: Write the equation of the line with slope 2 that passes through the point (1, −3). y = ax + b. This might be able to help you :) http://www.themathpage.com/alg/slope-of-a-line-2.htm

OpenStudy (anonymous):

do the answer c

OpenStudy (anonymous):

so

OpenStudy (anonymous):

I believe so. What do you think?

OpenStudy (anonymous):

i think it is

OpenStudy (anonymous):

Then I guess that's your answer :D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome. Tag me if you need any more help. :)

OpenStudy (anonymous):

Select the equation of the line in slope-intercept form: 4x + 9y = –36